# AI Copilots for Reproducibility in Science: A Case Study

### v3

@@ Abstract
Open science initiatives seek to make research outputs more transparent, accessible, and reusable, but ensuring that published findings can be independently reproduced remains a persistent challenge. This paper introduces OpenPub, an AI-powered platform that supports researchers, reviewers, and readers through a suite of modular copilots focused on key open science tasks. In this work, we present the Reproducibility Copilot, which analyzes manuscripts, code, and supplementary materials to generate structured Jupyter Notebooks and recommendations aimed at facilitating computational, or “rote”, reproducibility. We conducted feasibility tests using previously studied research papers with known reproducibility benchmarks. Results indicate that OpenPub can substantially reduce reproduction time - from over 30 hours to about 1 hour - while achieving high coverage of figures, tables, and results suitable for computational reproduction. The system systematically detects barriers to reproducibility, including missing hyperparameters, undocumented preprocessing steps, and incomplete or inaccessible datasets. While preliminary, these findings suggest that AI-driven tools can meaningfully reduce the burden of reproducibility efforts and contribute to more transparent and verifiable scientific communication. The modular copilot architecture also provides a foundation for extending AI assistance to additional open science objectives beyond reproducibility.
@@ Introduction
The scientific enterprise is a highly collaborative effort, with research conducted by groups of scientists working in separate labs that are often scattered around the world. As such, scientific progress is critically dependent on the accurate, truthful communication and curation of technical results. The scientific publication process, in particular the journal publication process, has been the primary means of communicating technical results for hundreds of years.
The desire to share and reuse scientific data and software has grown in recent years as funders (e.g., National Science Foundation [1, 2, 3]) and publishers (e.g., American Geophysical Union [4]) have introduced open science policies that emphasize reproducibility, and as scientists increasingly recognize the benefits of open science [5, 6]. Consequently, the last decade has seen a proliferation of frameworks (e.g., electronic notebooks and containers) to promote open source resources, collaboration among scientists, and sharing of science products.
The advent of the Web also ushered in the Open Access Movement, lowering the barrier of subscription charges, and the Web environment has also encouraged experimentation with a wide variety of approaches to reviewing and feedback. More recently, the concept of “Open Science,” which broadly refers to the open and transparent sharing of data, physical samples, software, models, knowledge, etc., has also become a rallying point for the scientific community [7].
As discussed by Gil et al. (2016), the practice of open science can require additional time and work by researchers, such as documenting code and data so that analyses can be reproduced [8]. However, AI can now have a high impact in helping researchers in this endeavor. Indeed, the striking new advances in AI over the past few years, largely driven by the success of large language models (LLMs), have created new opportunities to improve how science is conducted and potentially revolutionize current paradigms for reviewing, curating, publishing, distributing, and sharing new technical results.
In this paper, we present a case study that shows how large the impact of AI can have on making science more open. To make that possible, we are developing OpenPub, an AI-enabled set of copilots for helping the different actors (authors, readers, reviewers, publishers, etc.) in the open science enterprise. In this study, we focus on a copilot that helps authors make their research more reproducible and that helps readers to reproduce the authors’ research. We developed an initial version of this copilot, and demonstrated the feasibility of the approach. For instance, we revisited a study on reproducibility by Gundersen et al. [9], and showed that for a paper that previously took 30 hours to reproduce, our copilot could produce recommendations which, if followed, would reduce the time to reproduce the paper to one hour.
@@
In considering what problems the Reproducibility Copilot should tackle, we were influenced by the analysis of Gundersen et al. (2025), who analyzed the impact of different issues related to making science more open [9]. Their categorization included impactful issues, such as missing hyperparameter values for reproducing experiments, missing or undocumented code components, and incomplete data references. Using Gundersen et al. analysis as a guide, we designed a copilot capable of detecting and providing recommendations to fix a spectrum of problems.
Figure 1 provides a high-level overview of the architecture of the system, which is designed to support both the authors and the readers in a coordinated back-and-forth manner. This dual role design allows the system to anticipate the reader’s perspective, identify elements that can be difficult to interpret or reproduce, and generate a targeted output for the reader (e.g., a Jupyter Notebook to guide reproduction) while also producing actionable recommendations for the author (e.g., comments in the manuscript and code annotations).
Fig 1: System architecture of OpenPub, our AI-driven Reproducibility Copilot, illustrating its dual-path design that simultaneously supports authors and readers.
@@
The generated Notebook, shown in Figure 2, begins with a clear structure to reproduce the figures and tables of the study. By organizing the reader’s process and clarifying where input is needed, the Notebook serves as a scaffold rather than a replacement for the author’s work. In Figure 3, we show an example cell that corresponds to a figure from Rodriguez and Laio (2014) [10]. The system has added a placeholder to indicate that the code to generate the plot is missing, and uses this signal to later generate a recommendation to the author to supply the relevant code.
Fig 2: Opening section of the Jupyter Notebook automatically generated by the system to guide readers in reproducing the experiments from Rodriguez and Laio (2014) [10].
Fig 3: Excerpt from the system-generated Notebook for reproducing Figure 4 from Rodriguez and Laio (2014) [10]. Since the original code for this figure was not provided, the system inserts minimal placeholder content to signal the missing implementation.
@@
    Code Checker: Examines code to detect missing snippets essential for experiment replication.
@@
Each of these modules is implemented using GPT-4o [11] and a 2-step prompt strategy. For the first step, a prompt was designed to list candidates in each specific module. For instance, in the case of the Hyperparameter Checker, this first step consisted in listing all the hyperparameters mentioned in the manuscript and code. The second step then uses a specific prompt to reduce the list according to the issue at hand and provides some recommendations to the authors. In the case of the Hyperparameter Checker, this consisted in keeping the hyperparameters for which some values are not clearly provided for some experiments. Given the non-deterministic nature of GPT-4o, these checks were performed five times, and another prompt was used with GPT-4o to consolidate these results by taking the union of the final lists produced.
Once all module outputs are consolidated, the final list produced by each module is used by the system to annotate the manuscript PDF with targeted highlights and margin comments, and inserts inline feedback within code files. As the number of comments can be high, depending on the reproducibility issues in the paper, the corresponding feedback comments were deliberately kept minimal to avoid overwhelming the authors.
@@ Results
A crucial result of our case study is the demonstration of a significant time reduction in reproducing scientific research. To demonstrate the practical impact of the system, we selected a paper by Rodriguez and Laio (2014) [10], which had previously been analyzed in detail by Gundersen et al. (2025) [9]. This paper, which introduced a new machine learning method for clustering, was chosen specifically because it was one of the studies successfully reproduced by Gundersen et al., confirming that replication of the results was possible. Furthermore, the reported reproduction time of 33 hours by Gundersen et al. was below the 40-hour cutoff used in their study. Since this reproduction effort did not reach the cutoff limit, the 33-hour figure reflects the complete time it took to fully reproduce the study, making it a reliable benchmark against which we could measure our system’s efficiency.
@@
Although it took a researcher 33 hours to reproduce the experiments in the Gundersen et al. study, using our copilot system, we found that users could complete the task in a relatively short time. Specifically, we asked two individuals (who were unfamiliar with the reproduced paper, its methodology, and its data) to independently play the role of readers and use our copilot to replicate the study. The first user, a researcher with a Ph.D. in data science, completed the task in about 1 hour. The second, a somewhat less experienced user (a college student in computer science with substantial AI experience), also achieved similar performance (i.e., about 1 hour to reproduce the experiments), suggesting that the system makes reproducibility more accessible, not just more efficient.
@@
These tests were carried out under the assumption that the author is fully compliant with the recommendations generated by our tool, allowing us to evaluate the effectiveness of the system in an ideal usage scenario. What this means is that the copilot generated suggestions for how to modify the code and documentation, then we (members of our project team) played the role of the authors, modifying and/or adding code, data, and documentation according to the system’s suggestions. To emulate the role of the authors, we reused and adapted publicly available code: the supplementary material provided by Rodriguez and Laio (2014) [10] and the implementation from Gundersen et al. (2018) (available at https://github.com/AIReproducibility2018/ClusteringByFastSearchAndFindOfDensityPeaks). Once the copilot was satisfied, the paper and supplementary material was made available to the user we recruited for our reproducibility task. In practice, for the paper by Rodriguez and Laio (2014) [10], working with our system produced the data for each experiment and a Notebook containing functions to load each file in the right format for the paper’s method, the specific hyperparameter values for each of the experiments, and a short explanation on how to use the paper’s methods and analyze each result. These, together, were the missing pieces that made the reproduction of the paper take more than 30 hours in the study by Gundersen et al.
@@
To evaluate content coverage, we conducted a separate experiment in which we compared the Jupyter Notebook generated by our copilot with one that was manually created. This experiment focused on an Earth Sciences study, in part to test our prototype on another type of subject. Specifically, the study used Jupyter Notebooks [12] that were developed by Dr. Deborah Khider (one of the coauthors of this paper) to help teach the research behind a study by DeLong et al. (2011) [13]. The goal of these pedagogical Notebooks was not full reproducibility, but rather a guide for early-career scientists to learn to use data science libraries such as scikit-learn to perform a calibration task. While the original study focused on the robustness of the proxy signal in space and among different species of corals, the pedagogical Notebooks illustrate the use of Frequentist and Bayesian approaches in performing this calibration task. In this case, to mimic the original authors following the system’s suggestions, we adapted the code of Dr. Deborah Khider (available at https://khider.github.io/dry-tortugas-calibration-fun).
When comparing the Notebook produced by OpenPub with the pedagogical ones from Dr. Khider, we found that OpenPub provided a wide coverage of figures and tables, capturing not only the core figures and tables highlighted in the educational Notebooks, but also additional elements of the paper that are essential for full reproduction. As shown in Table 1, OpenPub covered all key figures and tables included in the educational Notebooks (green rows), as well as several additional visualizations and results that are necessary from a reproducibility point of view, but were excluded in the pedagogical version (white rows). The only omission from the coverage of OpenPub was a figure (Figure 8 in DeLong et al. (2011) [13]), which was also not prioritized in the teaching-oriented Notebooks (orange row). This comprehensive coverage demonstrates that such an AI system does not merely replicate what educators consider useful for instruction: it successfully identifies and scaffolds the full scope of reproducible content in a research paper, showing that this kind of tool can be quite complete when used for scientific validation.
Table 1: Comparative analysis of reproducibility coverage between OpenPub’s generated Notebook and Dr. Khider’s educational Notebooks [12] for the DeLong et al. (2011) study [13]. Green rows indicate figures and tables covered by both systems, orange rows mark a figure not covered by OpenPub but also excluded from the educational Notebooks, and white rows highlight additional figures and tables captured by OpenPub but omitted from the educational Notebooks.
Figures/Tables from DeLong et al. (2011) OpenPub’s Notebook Educational Notebooks
@@
Figure 7 $\surd$ $\surd$
Figure 8 $\times$ $\times$
@@
Table 3 $\surd$ $\surd$
@@ Discussion
During our research, we identified several challenges related to the use of AI to enhance reproducibility.
@@
Reproducibility can take many forms, such as computational reproducibility [14, 15, 16], empirical reproducibility [17], and outcome, analysis and interpretation reproducibility [18]. Throughout the use case in this paper, we focused on what can be called “rote reproducibility”, which can be defined as regenerating figures and tables using the same code and data. This type of reproducibility is related to a more complete “scientific reproducibility”, where the reconstruction of the reasoning and conclusions behind the research is sought. However, achieving complete scientific reproducibility is substantially more difficult as it requires access to the logic, rationale, and underlying assumptions that may not be explicitly stated in a paper.
In this case study, only preliminary steps were taken towards the challenging issues of scientific reproducibility. For example, we inferred analytical workflows from the manuscript and added preliminary instructional steps in the generated Notebook to guide readers in evaluating the plausibility of reproduced results. However, making it possible for the system to have a full understanding of the science behind the paper is a challenge to tackle in future work.
@@
A related complication comes from differences in the scientific background of the readers. What may be intuitive to an expert may be opaque to a student or newcomer. The tailoring of the guide based on the reader’s expertise is essential to promote scientific understanding. In future work, we plan to automatically construct semantic representations of the author analyses, enabling AI copilots to generate reader-specific guidance. This would not only support reproducibility, but also accelerate comprehension and learning for reviewers and readers alike. Indeed, our observations during our study suggest that simply understanding a paper often requires as much effort as reproducing the study using our Notebook.
@@
Reproducing a study often (but not always) involves executing the code provided by the author. In some cases, there are bugs in that code, and in many cases, some parts of the code are missing, or at least, there is code that the reader needs to write to conveniently reproduce the work. Though in many cases we found that LLMs could do a fairly good job of automatically generating such code, we felt that this would be a distraction in this case study (especially if the generated code contained subtle errors), and we decided to rely on the authors to enter the code in the Notebook (who often actually has the code somewhere). In future work, we plan to reconsider how OpenPub can make use of the rapidly improving capabilities of modern LLMs to help with bug-free code generation as part of the reproducibility process.
@@
This study presents a feasibility study rather than a large-scale evaluation. Furthermore, our experiments assumed full compliance by the original authors with the system, which may not always be realistic in practice. Finally, the use of large language models introduces some variability due to non-deterministic outputs. These limitations underscore the preliminary nature of our findings, but also highlight clear directions for future work.

### v4

@@ Abstract
Open science initiatives seek to make research outputs more transparent, accessible, and reusable, but ensuring that published findings can be independently reproduced remains a persistent challenge. In this paper we describe an AI-driven ”Reproducibility Copilot” that analyzes manuscripts, code, and supplementary materials to generate structured Jupyter Notebooks and recommendations aimed at facilitating computational, or “rote”, reproducibility. Our initial results suggest that the copilot has the potential to substantially reduce reproduction time (in one case from over 30 hours to about 1 hour) while achieving high coverage of figures, tables, and results suitable for computational reproduction. The system systematically detects barriers to reproducibility, including missing values for hyperparameters, undocumented preprocessing steps, and incomplete or inaccessible datasets. Although preliminary, these findings suggest that AI tools can meaningfully reduce the burden of reproducibility efforts and contribute to more transparent and verifiable scientific communication.
Codehttps://bitbucket.org/inferlink/reproducibility˙copilot
Notebookshttps://zenodo.org/records/17162894
@@ Introduction
Open science has emerged as a transformative movement aimed at improving the transparency, accessibility, and reliability of scientific knowledge (OpenScienceBook). Its importance stems from the recognition that science advances most effectively when findings, methods, and data are openly shared and critically examined. Open science strengthens the scientific enterprise by making it easier for readers to transparently investigate published results and build on prior work, thereby accelerating discovery and enhancing the reliability of scientific claims.
Among the pillars of open science, reproducibility is particularly central. Reproducibility ensures that scientific results are not isolated artifacts but can be independently verified and built upon by others. Without it, trust in scientific claims is weakened and progress becomes difficult to assess or scale. Although some researchers have long championed open practices (boai2002; mckiernan2016open; lowndes2017our), the past decade has seen a significant acceleration, driven in part by mandates from funders and publishers (nugroho2015comparison; vandeneyden2016survey; zuiderwijk2014open; doerr2019giving), and by investments in open science infrastructure - such as curated data repositories, community standards, and tools such as electronic notebooks and containers - alongside the automation workflows that streamline the use of these tools by scientists.
As discussed by GeosciencePaperOfTheFuture, the practice of open science can require additional time and work by researchers, such as documenting code and data so that analyses can be reproduced. However, artificial intelligence (AI), as well as electronic notebooks and containers, can now help researchers in this endeavor. Indeed, the striking new advances in AI in recent years, including in large language models (LLMs), have created new opportunities to assist researchers and potentially revolutionize current paradigms, including reviewing, curating, publishing, distributing and sharing new technical results (see, for instance, the work of Paper2Agent and the new Agents4Science conference [^1] https://agents4science.stanford.edu/). To take advantage of these capabilities, we have been developing an AI-copilot framework called OpenPub to help authors and readers, but also reviewers and publishers, in the open science enterprise.
In this paper, we present a case study on an AI copilot that 1) helps authors make their research more reproducible and 2) helps readers to reproduce the authors’ research. We developed an initial version of this Reproducibility Copilot and demonstrated the feasibility of the approach. For instance, we revisited a study on reproducibility by gundersen2025unreasonableeffectiveness, and showed that for a paper that previously took 30 hours to reproduce, our copilot could produce recommendations which, if followed, would reduce the time to reproduce the paper to approximately one hour.
@@
In considering what problems the Reproducibility Copilot should tackle, we were influenced by the analysis of gundersen2025unreasonableeffectiveness, who analyzed the impact of different issues related to making science more open. Their categorization included impactful issues, such as missing hyperparameter values for reproducing experiments, missing or undocumented code components, and incomplete data references. Using gundersen2025unreasonableeffectiveness analysis as a guide, we designed a copilot capable of detecting and providing recommendations to fix a spectrum of problems.
Figure 1 provides a high-level overview of the architecture of the system, which is designed to support both authors and readers. This dual-role design allows the system to anticipate the reader’s perspective, identify elements that can be difficult to interpret or reproduce, and generate a targeted output for the reader (e.g., a Jupyter Notebook to guide reproduction) while also producing actionable recommendations for the author (e.g., comments in the manuscript and code annotations). We observed that the iterative process of our system typically converges within one or two iterations. This rapid stabilization occurs because the LLM handles the majority of the interpretative workload upfront, while simultaneously prompting the author to supply the bulk of any missing context during the initial interaction. Consequently, subsequent iterations yield diminishing marginal gains.
Figure 1: System architecture of our AI-driven Reproducibility Copilot, illustrating its dual-role design that simultaneously supports authors and readers. While code and additional docs enhance results, they are optional - the system can function with the manuscript alone, generating PDF highlights to identify missing reproducibility information.
@@
The generated Notebook, shown in Figure 2, focuses on the reproduction of the figures and tables of the study. [^2] We have observed that most results tend to appear in figures or tables. So, for simplicity, results that are not mentioned in any table or figure are effectively ignored in the current implementation. By organizing the reader’s process and clarifying where input is needed, the Notebook serves as a scaffold rather than a replacement for the author’s work. In Figure 3, we show an example cell that corresponds to a figure from rodriguez2014clustering. The system has added a placeholder to indicate that the code to generate the plot is missing (see “# Add code to generate and display the results”), and uses this signal to later generate a recommendation to the author to supply the relevant code.
Figure 2: Opening section of the Jupyter Notebook automatically generated by the system to guide readers in reproducing the experiments from rodriguez2014clustering.
Figure 3: Excerpt from the system-generated Notebook for reproducing Figure 4 from rodriguez2014clustering. Since the original code for this figure was not provided, the system inserts minimal placeholder content to signal the missing implementation.
@@
    Code Checker: Examines the code to detect missing snippets essential for experiment replication.
@@
Each of these modules was implemented using GPT-4o (GPT-4o) and a 2-step prompt strategy. For the first step, a prompt was designed to list candidates in each specific module. For instance, in the case of the Hyperparameter Checker, this first step consisted in listing all the hyperparameters mentioned in the manuscript and code. The second step then uses a specific prompt to reduce the list according to the issue at hand and provides some recommendations to the authors. In the case of the Hyperparameter Checker, this consisted in keeping the hyperparameters for which some values are not clearly provided for some experiments. Given the non-deterministic nature of GPT-4o, these checks were performed five times, and another prompt was used with GPT-4o to consolidate these results by taking the union of the final lists produced.
Figure 4: Example of a highlight produced by the copilot on the manuscript PDF of rodriguez2014clustering. The associated comment suggests including data to ensure the reproducibility of the authors’ results.
Once all module outputs are consolidated, the final list produced by each module is used by the system to annotate the manuscript PDF with targeted highlights and margin comments (see, e.g., Figure 4), and inserts inline feedback within code files. As the number of comments can be high, depending on the reproducibility issues in the paper, the corresponding feedback comments were deliberately kept minimal to avoid overwhelming the authors.
@@ Results
A crucial result of our case study is the demonstration of a significant time reduction in reproducing scientific research. To demonstrate the practical impact of the system, we selected a paper by rodriguez2014clustering, which had previously been analyzed in detail by gundersen2025unreasonableeffectiveness. This paper, which introduced a new machine learning method for clustering, was chosen specifically because it was one of the studies successfully reproduced by Gundersen et al., confirming that replication of the results was possible. Furthermore, the reported reproduction time of 33 hours by Gundersen et al. was below the 40-hour cutoff used in their study. Since this reproduction effort did not reach the cutoff limit, the 33-hour figure reflects the complete time it took to fully reproduce the study, making it a reliable benchmark against which we could measure our system’s efficiency.
@@
Although it took a researcher 33 hours to reproduce the experiments in the Gundersen et al. study, using our copilot system we found that users could complete the task in a relatively short time. Specifically, we asked two individuals (who were unfamiliar with the reproduced paper, its methodology, and its data) to independently play the role of readers and use our copilot to replicate the study. The first user, a researcher with a Ph.D. in data science, completed the task in about 1 hour. The second, a somewhat less experienced user (a college student in computer science with substantial AI experience), also achieved similar performance (i.e., about 1 hour to reproduce the experiments), suggesting that the system makes reproducibility more accessible, not just more efficient.
@@
These tests were carried out under the assumption that the author is fully compliant with the recommendations generated by our tool, allowing us to evaluate the effectiveness of the system in an ideal usage scenario. What this means is that the copilot generated suggestions for how to modify the code and documentation, then we (members of our project team) played the role of the authors, modifying and/or adding code, data, and documentation according to the system’s suggestions. To emulate the role of the authors, we reused and adapted publicly available code: the supplementary material provided by rodriguez2014clustering and the implementation from gundersen2025unreasonableeffectiveness 333available at https://github.com/AIReproducibility2018/ClusteringByFastSearchAndFindOfDensityPeaks. Once the copilot was satisfied, the paper and supplementary material were made available to the user we recruited for our reproducibility task. In practice, for the paper by rodriguez2014clustering, working with our system produced the data for each experiment and a Notebook containing functions to load each file in the right format, specific hyperparameter values, and usage explanations. The system facilitated this by highlighting relevant sections of the original code and manuscript, prompting the author to clarify ambiguities and populate the Notebook with the correct context. Because the author was prompted to address these details, the recruited users were spared the manual reverse-engineering that made the reproduction of the paper take more than 30 hours in the study by gundersen2025unreasonableeffectiveness.
@@
To evaluate content coverage, we conducted a separate experiment in which we compared the Jupyter Notebook generated by our copilot with one that was manually created. This experiment focused on an Earth Sciences study, in part to test our prototype on another type of subject. Specifically, the study used Jupyter Notebooks that were developed by Khider_Coral_Sr_Ca_Calibration_2024 (one of the coauthors of this paper) to help teach the research behind a study by Tortugas. The goal of these pedagogical Notebooks was not full reproducibility, but rather a guide for early-career scientists to learn to use data science libraries such as scikit-learn to perform a calibration task. While the original study focused on the robustness of the proxy signal in space and among different species of corals, the pedagogical Notebooks illustrate the use of Frequentist and Bayesian approaches in performing this calibration task. In this case, to mimic the original authors following the system’s suggestions, we adapted the code of Dr. Deborah Khider 444available at https://khider.github.io/dry-tortugas-calibration-fun.
When comparing the Notebook produced by our copilot with the pedagogical ones from Dr. Khider, we found that our copilot provided a wide coverage of figures and tables, capturing not only the core figures and tables highlighted in the educational Notebooks, but also additional elements of the paper that are essential for full reproduction. As shown in Table 1, our copilot covered all key figures and tables included in the educational Notebooks (green rows), as well as several additional visualizations and results that are necessary from a reproducibility point of view, but were excluded in the pedagogical version (white rows). The only omission from the coverage of our copilot was a figure (Figure 8 in Tortugas), which was also not prioritized in the teaching-oriented Notebooks (orange row). This comprehensive coverage demonstrates that such an AI system does not merely replicate what educators consider useful for instruction: it successfully identifies and scaffolds the full scope of reproducible content in a research paper, showing that this kind of tool can be quite complete when used for scientific validation.
Figures/Tables from Tortugas Our Copilot’s Notebook Educational Notebooks
@@
[…]
@@
[…]
Table 1: Comparative analysis of reproducibility coverage between the Notebook generated by our copilot and Dr. Khider’s educational Notebooks (Khider_Coral_Sr_Ca_Calibration_2024) for the Tortugas study. Green rows indicate figures and tables covered by both systems, orange rows mark a figure not covered by our copilot but also excluded from the educational Notebooks, and white rows highlight additional figures and tables captured by our copilot but omitted from the educational Notebooks.
@@ Discussion
During our research, we identified several future challenges related to the use of AI to enhance reproducibility.
@@
Reproducibility can take many forms, such as computational reproducibility (freire2012computational; piccolo2016tools; gruning2018practical), empirical reproducibility (stodden2013resolving), and outcome, analysis and interpretation reproducibility (gundersen2022sourcesOfirreproducibility). Throughout the use case in this paper, we focused on what can be called “rote reproducibility”, which we define as regenerating figures and tables using the same code and data. This type of reproducibility is only a small part of full “scientific reproducibility”, where the reconstruction of the reasoning and conclusions behind the research is sought. However, achieving complete scientific reproducibility is substantially more difficult as it requires access to the logic, rationale, and underlying assumptions that may not be explicitly stated in a paper.
In this case study, only preliminary steps were taken towards the challenging issues of scientific reproducibility. For example, we inferred analytical workflows from the manuscript and added preliminary instructional steps in the generated Notebook to guide readers in evaluating the plausibility of reproduced results. Developing a copilot with a full understanding of the science behind the paper is a challenge to tackle in future work and would require knowledge beyond analysis.
@@
A related complication comes from differences in the scientific background of the readers. What may be intuitive to an expert may be opaque to a student or newcomer. The tailoring of the guide based on the reader’s expertise is essential to promote scientific understanding. In future work, we plan to automatically construct semantic representations of the author analyses, enabling AI copilots to generate reader-specific guidance. This would not only support reproducibility, but also accelerate comprehension and learning for reviewers and readers alike. Indeed, our observations during our study suggest that simply understanding a paper often requires as much effort as reproducing the study using the Notebook produced by our copilot.
@@
Reproducing a study often (but not always) involves executing the code provided by the author. In some cases, there are bugs in that code, and in many cases, some parts of the code are missing, or at least, there is code that the reader needs to write to conveniently reproduce the work. Though in many cases we found that LLMs could do a fairly good job of automatically generating such code, we felt that this would be a distraction in this case study (especially if the generated code contained subtle errors), and we decided to rely on the authors to enter the code in the Notebook (who often actually has the code somewhere). In future work, we plan to reconsider how our copilot can make use of the rapidly improving capabilities of modern LLMs to help with bug-free code generation as part of the reproducibility process.
Multi-Turn Interactions
Similar to coding agents like Codex of OpenAI (Codex) and Claude Code of Anthropic (Claude_Code), our Reproducibility Copilot could be enhanced with multi-turn interactions. In its current form, the copilot identifies potential improvements, allowing authors to revise their material for re-evaluation. However, challenges may arise if authors misinterpret the copilot’s recommendations or require specific guidance on implementation (e.g., how to properly archive data). To address this, the system would benefit from a conversational interface that assists authors in resolving these reproducibility barriers interactively.
@@
This study presents a feasibility study rather than a large-scale evaluation. A comprehensive evaluation could involve a controlled study where multiple PhD students reproduce papers in their respective fields, comparing the performance of those using the copilot against a control group working manually. Furthermore, our experiments assumed full compliance by the original authors with the system, which may not always be realistic in practice. Finally, the use of large language models introduces some variability due to non-deterministic outputs. These limitations underscore the preliminary nature of our findings, but also highlight clear directions for future work.